-2(4-3x)+7=6(x+1)
solve for x
: -2(4-3x)+7=6(x+1)
solve for x

Answer by checkley77(3654)

-2(4-3x)+7=6(x+1)
-8+6x+7=6x+6
6x-6x=6-7+8
0=7 no solution for x.